OpenStudy (anonymous):

Find an equation of the tangent line to the graph of "f" at the given point. f(x)=x²+3, (1,4)

OpenStudy (anonymous):

do you know what the derivative of \[f(x)=x^2+3\] is?

OpenStudy (anonymous):

No

myininaya (myininaya):

do you know the power rule or the constant rule?

OpenStudy (anonymous):

Yes, I do.

myininaya (myininaya):

Then use them to find the derivative.

myininaya (myininaya):

Power Rule \[(x^n)'=nx^{n-1} \] Constant Rule (C is a constant) \[(C)'=0\]

myininaya (myininaya):

\[f'(x)=(x^2)'+(3)'=?\]

OpenStudy (anonymous):

2x?

myininaya (myininaya):

yep

myininaya (myininaya):

f'(1) will be the slope remember f'(x)=2x so what is f'(1)=?

OpenStudy (anonymous):

2

myininaya (myininaya):

so we know this about the tangent line y=2x+b so we need one more thing, the y-intercept

myininaya (myininaya):

You know a point on this line (1,4) Plug it in to find b, the y-intercept.

myininaya (myininaya):

\[4=2(1)+b\] I replaced x with 1 and y with 4.

myininaya (myininaya):

find b

OpenStudy (anonymous):

b=2

myininaya (myininaya):

so the tangent line is y=mx+b where m=2 and b=2 y=2x+2

OpenStudy (anonymous):

alright, thank you

OpenStudy (anonymous):

A plot is attached.